OpenStudy (nopeforce):

Which ordered pair is a solution to the equation 3x + 5y = 9? A. (–2, –3) B. (2, 3) C. (2, –3) D. (–2, 3) I think c

rebeccaxhawaii (rebeccaxhawaii):

3(2)+5(-3)=6-15=-9

rebeccaxhawaii (rebeccaxhawaii):

so no not C

OpenStudy (nopeforce):

but its positive 9?

OpenStudy (anonymous):

D.

OpenStudy (mathmale):

Welcome to OpenStudy! Since you are given only ONE equation in TWO unknowns, you cannot find x and y without further info. However, you can substitute the x- and y-coordinates of each given point into the given equation to see whether the equation is or is not true. If not true, keep going until you identify the point whose coordinates do satisfy the given equation.

rebeccaxhawaii (rebeccaxhawaii):

it would be D 3(-2)+5(3)=-6+15=9

rebeccaxhawaii (rebeccaxhawaii):

C would come out to -9 which is not what we are looking for

OpenStudy (mathmale):

It's EQUALITY that is important here. The two sides of the equation must be EQUAL. If they are not, the equation is not true. In such case, move on to the next point and repeat the process.

OpenStudy (nopeforce):

thanks, mathmale really cleared it out for me

OpenStudy (mathmale):

Delighted to hear that.

OpenStudy (nopeforce):

how do i give metals!?

OpenStudy (mathmale):

Thank you for wanting to do that. Just click on the Best Response button. Appreciated.

OpenStudy (nopeforce):

it says ive already done that sorry :(

OpenStudy (mathmale):

It's the thought that counts! Good going, best wishes to you.